OpenStudy (anonymous):

Limit of lnx-ln(x^2+1) as x goes to infinity

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}[\ln(x)-\ln(x^2+1)] \\ \lim_{x \rightarrow \infty}\ln(\frac{x}{x^2+1})\]

OpenStudy (freckles):

where is the inside going

OpenStudy (freckles):

x/(x^2+1) goes to what as x->infty?

OpenStudy (anonymous):

oh duh

OpenStudy (anonymous):

then just use l'hopital

OpenStudy (anonymous):

completely forgot about that limit property

OpenStudy (freckles):

\[u=\frac{x}{x^2+1} \\ u \rightarrow\text{ to what ? } \text{ as } x \rightarrow \infty \] \[\lim_{x \rightarrow \infty}\frac{x}{x^2+1}\] you don't need l'hospital for this but you can

OpenStudy (anonymous):

wait wouldnt the limit go to 0?

OpenStudy (freckles):

\[\lim_{u \rightarrow 0}\ln(u)=?\]

OpenStudy (freckles):

I put u->0 as x->infty since as x->infty u=x/(x^2+1)->0

OpenStudy (anonymous):

So then then lim as x goes to infinity IS 0

OpenStudy (freckles):

well that is for the inside

OpenStudy (freckles):

it isn't really cool to say this but what is ln(0)

OpenStudy (anonymous):

isnt the lim of ln(x/x^2+1) equal to ln of lim of (x/x^2+1)?

OpenStudy (freckles):

yeah

OpenStudy (freckles):

sorta

OpenStudy (anonymous):

well ln(0) goes towards negative infinity

OpenStudy (freckles):

right

OpenStudy (freckles):

I liked doing a substitution

OpenStudy (anonymous):

ok i'll try it with a substitution but ive never done it that way before

OpenStudy (freckles):

it makes more sense to me \[\text{ \let } u=\frac{x}{x^2+1} \\ \text{ since } \lim_{x \rightarrow \infty} u =\lim_{x \rightarrow \infty}\frac{x}{x^2+1}=0 \\ \text{ so we have } \lim_{u \rightarrow 0} \ln(u)=-\infty \\ \text{ so we are saying } \\ \lim_{x \rightarrow \infty}\ln(\frac{x}{x^2+1})=\lim_{u \rightarrow 0}\ln(u)=-\infty \] But this does make sense/

OpenStudy (freckles):

that was suppose to end with a ?

OpenStudy (anonymous):

-infinity